OpenStudy (anonymous):

On a 432 mile road trip, you determine that you stopped at a gas station and then a rest stop in between the starting and ending destination of your trip. If the gas station was located halfway between your start and endpoint and the rest stop was located halfway in between the gas station and your endpoint what was the distance from the rest stop to the ending destination?

OpenStudy (theloshua):

easy first split the distance in half

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }~of 432 miles=?\]

OpenStudy (theloshua):

432/2 = 216

OpenStudy (theloshua):

then split that in half 216/2 = 108

OpenStudy (theloshua):

which should be your answer

OpenStudy (anonymous):

Yup, I agree

OpenStudy (theloshua):

*pats myself on the back* lol

OpenStudy (anonymous):

Thanks

OpenStudy (theloshua):

no probs :D